OpenStudy (anonymous):

Two coins are flipped 1,000 times with the following frequencies 2 tails 210 1 tail 480 0 tail 310. Empirical and theoretical probability

OpenStudy (lyrae):

Empirical 2 tail: 210/1000 = 0.21 1 tail: 480/1000 = 0.48 0 tail: 310/1000 = 0.31 Theoretical 2 coins gives four possible outcomes with equal probabillity every time you throw: Coin 1 Coin 2 Probabillity 1: tail tail 1/4 2: head tail 1/4 3: tail head 1/4 4: head head 1/4 In this particular case, 2 and 3 are the same (1 tail) so their probabillity is added 1/4 + 1/4 = 2/4 = 1/2 This gives us the final result: 2 tail: 1/4 = 0.25 1 tail: 1/2 = 0.50 0 tail 1/4 = 0.25

OpenStudy (anonymous):

Thank you so much!,

OpenStudy (anonymous):

a certain dry cereal is packaged on 2 different machines. machine a is set on aveage so that it fills the boxes with 24.5 oz of cereal and standard deviation of 0.4 oz. machine b is set so that it fills the boxes on average with 24 oz of cereal with a standard deviation of 0.2oz. describe which machine is more consistent and why